OpenStudy (anonymous):

Write a polynomial function of f of degree 5such that the end behavior of the graph of f is given by f(x)-->+infinity as x -->-infinity and f(x)-->-infinity as x -->+infinity. Then graph the function to verify your answer.

OpenStudy (jamesj):

One of these two will do it. Which? f(x) = x^5 g(x) = -x^5